OpenStudy (anonymous):

ok I got another one, I have been reading can' t find an example, I am sure it is easy but I just can't place it. 4y to the 3rd power over 8y 4y to the 3rd/8y

OpenStudy (lgbasallote):

\[\large \frac{4y^3}{8y} \implies \frac 48 \times \frac{y^3}{y} \implies \frac 12 \times y^{3-1}\] does that help?

OpenStudy (anonymous):

oh yes thank you so so much!!!

OpenStudy (lgbasallote):

welcome :D